OpenStudy (anonymous):

Evaluate the Integral: I just don't get how to format this into standard form :( Help much appreciated!

OpenStudy (anonymous):

\[\int\limits_{1}^{2}\frac{ (x-1^{3}) }{ x ^{2} }\]

OpenStudy (zehanz):

I think it should be: \[\int_{1}^{2}\dfrac{(x-1)^3}{x^2}dx\]

OpenStudy (anonymous):

aha oops you're right

OpenStudy (anonymous):

Expand the numerator, then divide.

OpenStudy (zehanz):

Expand the brackets -- ah @SithsAndGiggles said it!

OpenStudy (anonymous):

Ahhh okay... I was thinking of that but thought there would be a faster way. Thanks :)

OpenStudy (zehanz):

YW!